OpenStudy (anonymous):

simplify this expression help please! :( -20-(-5) * (-2^2)

OpenStudy (anonymous):

(-20+5) x 4 -15 x 4 -60

OpenStudy (anonymous):

thank you guys ! :)

OpenStudy (anonymous):

-20-(-5) * (-2^2) -20-(-5)*(-2^2) -20-(-5)*4 -20_(-20) _20+20 0 0

OpenStudy (anonymous):

i see what you did thank you !

OpenStudy (anonymous):

good job @cal360superman

OpenStudy (anonymous):

thanks, keep up the great work everyone :)

OpenStudy (nincompoop):

digest this: \[-2^2=-1(2)^2=-4\] meanwhile, \[(-2)^2=4\]

OpenStudy (anonymous):

ok